What is the probability that two random selected people are born : a) in February & b) in the different months ?

media%2F220%2F22026d3d-9eec-4f04-b6b1-e7
Answer #1

1.

A person can have birthday in any of the 12 months of the year.Another person can also have his/her birthday in any of the 12 months.Now,the total number of possible combinations is 12*12 , which equals 144.
This set of all possible combinations contains 1 identical month {i.e. february},which are our favorable outcomes.
Thus the probability =
(Number of favorable outcomes)/(Number of events in sample space)=
1/(12*12)=1/144

2. for this month other than in which first person is born = 11

so p(in different month) = 12.11/12.12 = 11/12